OpenStudy (anonymous):

The product of two numbers is 336, and their difference is 10. What are the two numbers?

OpenStudy (anonymous):

xy=336 x-y=10

OpenStudy (anonymous):

x=10+y

OpenStudy (agreene):

(-14,-24); (14,24)

OpenStudy (anonymous):

xy=336 (y+10)y=336 y^2+10y=336 y^2+10y-336=0 (y )(y )

OpenStudy (anonymous):

so the answer is 14 and 24?

OpenStudy (anonymous):

????????????

OpenStudy (agreene):

14 and 24 or -14 and -24 both pairs are solutions.